Starbucks has found that 26% of customers purchase more than one cup of coffee per day form their shops. Suppose you randomly choose 190 of today's Starbucks customers

Find the probability that no less than 30% of them purchased more than one cup of coffee from Starbucks.
Answer #1

Using central limit theorem,

P( < p) = P( Z < - p / sqrt( ( 1 - ) / n) )

So,

P( >= 0.30) = P( Z >= 0.30 - 0.26 / sqrt( 0.26 ( 1 - 0.26) / 190) )

= P(Z >= 1.257)

= 1 - P(Z < 1.257)

= 1 - 0.8956

= 0.1044